# Bridging the Gap Between LLMs and Human Intentions: Progresses and Challenges in Instruction Understanding, Intention Reasoning, and Reliable Generation

### v1

@@ 3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks Kim et al. (2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 4 C.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 4 C.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve the problem, researchers have started to experiment with cue engineering and relation alignment. Folkscope Yu et al. (2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko Lu et al. (2024) introduces a multimodal hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. ATC Deng et al. (2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
@@ 3.4 Intention Clarification Failure
Unlike humans, who reason based on experience, LLMs lack real-world common sense and thus struggle to infer complex contexts beyond their input data. Moreover, they often fail to maintain a consistent reasoning trajectory across long texts, complex contexts, or multi-turn conversations. When handling intricate intentions or sentiment shifts, LLMs may struggle to retain prior context, leading to errors in inferring implicit user needs. We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 4 D.1), when LLMs fails to understand the sarcastic intention of the user. 2) Ignores prior emotional context (Figure 4 D.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
To solve above problems, researchers have started to try to construct multi-domain datasets containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 Guo et al. (2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 Muennighoff et al. (2025) uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds ”wait” to guide in-depth reasoning and optimize the quality of the answer. SoulChat Chen et al. (2023) fine-tuned LLMs by constructing a dataset containing more than 2 million samples of multi-round empathic conversations. MoChat Mo et al. (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
4 Realiable Generation
Despite strong performance, LLMs struggle with output reliability. Trained on large corpora using maximum likelihood estimation, they generate deterministic responses. While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability. We analyze LLM response stability and alignment in depth.
@@ 4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al. (2019). If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 5 A.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 5 A.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs as follows.
Figure 5: Common generation issues in LLMs include: (A) Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (A.1) or produces incorrect contextual information despite having relevant knowledge (A.2). (B) Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (B.1) or provides responses that conflict with moral and ethical standards (B.2).
@@
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al. (2024). CalibrateMath Lin et al. (2022) assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
@@ 4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al. (2021). This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 5 B.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al. (2024); Dutta et al. (2024). 2) Conflicts with Moral/Ethical Standards (Figure 5 B.2). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al. (2024). To tackle the previously mentioned concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
@@
Table 1: A selection of widely used benchmark datasets for evaluating LLMs. The ’Type’ column categorizes the benchmarks based on task types, while ’Lang.’ indicates the language of the dataset, and ’Citation’ denotes the number of citations each benchmark has received.
Type Benchmark Year Lang. Description Citation
[…]
MT-BENCH-101 Bai et al. (2024) 2024 En Evaluation of multi-turn dialogue capability in LLMs. 39
LLM Reasoning BIPIA Yi et al. (2023) 2023 En Assesses LLMs vulnerability to hint injection attacks. 60
Miko Lu et al. (2024) 2024 En Evaluates multimodal LLMs in understanding social intention. 2
CONTRADOC Li et al. (2024a) 2023 En Detects self-contradictions in long documents. 7
CDCONV Zheng et al. (2022) 2022 Zh Detects contradictions in Chinese dialogues. 12
LLM Generation Open-LLM-Leaderboard Ye et al. (2024) 2024 En Assesses LLMs uncertainty in generated content. 19
ETHICS Dataset Hendrycks et al. (2021) 2020 En Evaluates LLMs on moral reasoning and ethics. 468
FACTOR Muhlgay et al. (2024) 2023 En Measures factual accuracy in LLM-generated text. 71
@@
To further align LLMs with human values and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al. (2022), DPO Zeng et al. (2024), and GRPO Shao et al. (2024), are very essential. As an extension of RLHF, RLAIF Lee et al. (2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
@@
In-context alignment leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. Lin et al. (2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. Huang et al. (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
@@ 5 Benchmark
This section covers benchmarks for LLMs in instruction understanding, reasoning, and generation (Table 1).
@@ 5.1 Benchmarking Instruction Understanding
Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues. Zhang et al. Zhang et al. (2024c) provides a benchmark consisting of 12 tasks, with an average data length exceeding 100K tokens, designed to evaluate LLMs’ ability to handle long contexts. The results show that the performance of LLMs significantly declines when dealing with long contexts, suggesting further improvements. MT-BENCH-101 Bai et al. (2024) constructs a three-layer hierarchical capability classification and 1,388 dialogue pairs across 13 tasks, providing a comprehensive evaluation of LLMs’ fine-grained capabilities in multi-turn dialogue. Experimental results show that commonly used techniques do not significantly enhance multi-turn capabilities.
@@ 5.2 Benchmarking LLM Reasoning
LLM Reasoning involves inferring user intentions by interpreting both explicit and implicit language cues. BIPIA Yi et al. (2023) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities. Miko Lu et al. (2024) assesses multimodal models in understanding social media user intentions. The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4. Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy. CONTRADOC Li et al. (2024a) is the first dataset for analyzing self-contradictions in long documents. Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning. CDCONV Zheng et al. (2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds. The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
@@ 5.3 Benchmarking LLM Generation
LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses. Open-LLM-Leaderboard Ye et al. (2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue. Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy. ETHICS Hendrycks et al. (2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being. The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain. FACTOR Muhlgay et al. (2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered. Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
@@ 6 Future Directions
This section summarizes ongoing challenges in instruction understanding, reasoning, and reliable generation with LLMs and outlines potential future research directions.
@@
Quantifying Uncertainty in LLMs.
As LLMs like LLama and ChatGPT have revolutionized content generation, ensuring the reliability of their outputs remains a critical challenge. Uncertainty quantification is a promising approach to address this, enabling models to provide a confidence assessment alongside their responses. Evidence learning Sensoy et al. (2018); Amini et al. (2020), an emerging method in uncertainty representation, offers a reliable approach for quantifying uncertainty directly from data. However, its application to LLMs is computationally intensive, and most existing work focuses on small-scale models. Future research should aim to optimize uncertainty quantification for large-scale models efficiently.
@@ 7 Conclusion
This paper analyzes LLMs’ performance in processing user instructions. Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations. Improvements through prompt engineering, knowledge retrieval, model expansion, fine-tuning, and RLHF have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability. We identify three key bottlenecks: instruction comprehension, reasoning ability, and reliable content generation. Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intent, to improve LLMs’ adaptability and reliability in real-world tasks.

### v2

@@ 3.3 Fuzzy Language Interpretation
When user instructions contain fuzzy terms (e.g., polysemy or vagueness), LLMs may select an incorrect interpretation from multiple possibilities, potentially leading to misleading responses. This phenomenon has been observed across multiple information-seeking tasks Kim et al. (2024), and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Self-defined problem (Figure 8 Q.1). When the user inputs content with fuzzy sentences, LLMs may choose to generate content based on the preferences of its own training data. 2) Select data based on fuzzy input (Figure 8 Q.2). In response to ambiguous user input, LLMs may select a default explanation without actively asking the user to clarify.
To solve this problem, researchers have started to parse the semantic information expressed by users through clue engineering. Folkscope Yu et al. (2023) proposed the FolkScope framework, which uses a large language model to analyze and discriminate users’ fuzzy purchasing intention. Miko Lu et al. (2024) introduces a hierarchical intention generation framework that interprets users’ posting behaviors by analyzing the fuzzy information they share on social platforms. ATC Deng et al. (2023) utilizes the Active Thinking Chain cueing scheme, which enhances the proactivity of a biglanguage model by adding goal-planning capabilities to the descriptive reasoning chain.
@@ 3.4 Intention Clarification Failure
Unlike humans, who reason based on experience, LLMs lack real-world common sense and thus struggle to infer complex contexts beyond their input data. Moreover, they often fail to maintain a consistent reasoning trajectory across long texts, complex contexts, or multi-turn conversations. When handling intricate intentions or sentiment shifts, LLMs may struggle to retain prior context, leading to errors in inferring implicit user needs. We categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fails to detect sarcasm (Figure 9 Q.1), when LLMs fails to understand the sarcastic intention of the user. 2) Ignores prior emotional context (Figure 9 Q.2), when LLMs focus only on the second half of the sentence and ignore the emotions in the previous round of dialog.
To solve above problems, researchers have started to try to construct multi-domain datasets Chen et al. (2023) containing implicit intentions to strengthen the ability of the LLMs to reason about complex intentions and user emotions in multi-round interaction scenarios. DeepSeek-R1 Guo et al. (2025) enhances its understanding of human intention through a structured process with two RL stages for refining reasoning patterns and aligning with human preferences. S1 Muennighoff et al. (2025) uses budget forcing to control the number of thinking tokens. The upper limit is terminated early by a delimiter, while the lower limit prohibits delimiters and adds "wait" to guide in-depth reasoning and optimize the quality of the answer. MoChat Mo et al. (2024) constructs multi-round dialogues for spatial localization by using joint grouping spatio-temporal localization.
Cat. Benchmark Year Lang. Num. Type Description
Instrution Understanding BotChat Duan et al. (2024) 2023 En&Zh 7658 M Multi-round dialogue eval. via simulated data
MINT Wang et al. (2024) 2023 En 29,307 M Tool use and feedback in multi-turn dialogue
MT-BENCH-101 Bai et al. (2024) 2024 En 1,388 M Multi-turn dialogue ability
$\infty${B}ench Zhang et al. (2024c) 2024 En&Zh 100k S Long-context handling
L-Eval An et al. (2024) 2024 En 2,000 S Evaluation of Long-Context Language Models
LongICLBench Li et al. (2025) 2025 En 2,100k S Long In-Context Learning
Intention Reasoning BIPIA Yi et al. (2025) 2023 En 712.5K S Vulnerability to hint injection
Miko Lu et al. (2024) 2024 En 10k S Multimodal social intent understanding
CONTRADOC Li et al. (2024a) 2023 En 891 S Self-contradiction in long docs
CDCONV Zheng et al. (2022) 2022 Zh 12K M Contradiction in Chinese dialogues
Relibale Dialog Generate Open-LLM-Leaderboard Ye et al. (2024) 2024 En 10K S Uncertainty in generation
ETHICS Hendrycks et al. (2021) 2020 En 130K S Moral reasoning
FACTOR Muhlgay et al. (2024) 2023 En 300 S Factuality in generated text
Table 1: A selection of widely used benchmark datasets for evaluating LLMs. Th“Cat.”: task stage; ’Lang.’: language of the benchmark, a; ’Num.’: data size; ’Type’: S=Single-round, M=Multi-round.
4 Reliable Dialog Generation
Despite strong performance, LLMs struggle with output reliability. Trained on large corpora using maximum likelihood estimation, they generate deterministic responses. While effective on familiar data, they often produce unstable or incomplete responses to unseen inputs, undermining reliability.
@@ 4.1 Response Stability
The knowledge acquired by the LLMs is generally determined in the pre-training stage and stored in a parameterized form. For data in a specific field, the current model is generally optimized by fine-tuning instructions so that it outputs what humans want Radford et al. (2019). If knowledge samples that the model has not seen are used in instruction-tuning, it will inevitably cause the model to give a definite response to unknown inputs, and there is a high probability that an answer will be fabricated. This is the over-confidence of the model that causes the model to output unreliable answers, and we categorize the causes of this problem according to the scenarios in which it occurs as follows: 1) Fabricated incorrect information (Figure 10 Q.1). When the model’s knowledge did not match the input question, it fabricated information that did not match the facts. 2) Incorrect context output (Figure 10 Q.2). When the model’s knowledge did match the input question, it output incorrect context information. To address these issues, researchers have explored uncertainty, which quantifies the credibility and stability of model outputs.
@@
Fine-tuning LLMs typically demands substantial computing resources and slow training; therefore, reducing computational overhead is crucial for improving efficiency. Researcher has proposed methods to evaluate the uncertainty of model outputs through external tools Liu et al. (2024). ConfidenceElicitation Xiong et al. (2024) is a new uncertainty measurement tool for large model outputs. CalibrateMath Lin et al. assesses uncertainty by requiring models to generate numerical answers with confidence levels, evaluating their reliability.
@@ 4.2 Alignment
Despite the impressive capabilities of large language models (LLMs), they have raised significant concerns regarding the unsafe or harmful content they may generate. LLMs are typically trained on vast datasets scraped from the internet, including inappropriate or harmful content Bender et al. (2021). This means that the models may inadvertently produce outputs misaligned with human values as follows: 1) Generation of Toxic Content (Figure 11 Q.1). LLMs may generate toxic content, such as hate speech or offensive comments, when asked to respond to sensitive topics Luong et al. (2024); Dutta et al. (2024). 2) Conflicts with Moral/Ethical Standards (Figure 11 Q.1). LLMs might produce outputs that conflict with moral or ethical standards, such as guiding illegal activities Ramezani and Xu (2023); Abdulhai et al. (2024). To tackle the above concerns regarding unsafe or harmful content produced by LLMs, researchers have focused on various stages:
@@
To further align LLMs with human and societal norms, reinforcement learning methods, such as RLHF and its advanced variants, such as PPO Ouyang et al. (2022), DPO Zeng et al. (2024), and GRPO Shao et al. (2024), are very essential. As an extension of RLHF, RLAIF Lee et al. (2024) leverages AI systems to assist in the feedback process, making evaluation and fine-tuning more scalable.
@@
It leverages the ability of LLMs to adapt their responses based on a few examples provided in the prompt. Lin et al. (2024) demonstrates that effective alignment can be achieved purely through ICL with just a few stylistic examples and a system prompt. Huang et al. (2024a) explored the effectiveness of different components of In-context alignment, and found that examples within the context are crucial for enhancing alignment capabilities.
5 Case Analysis
To demonstrate the challenges of existing LLMs in above stages, we provide cases analysis in Appendix B, and make statistical comparisons in Table 2. The results show that models with reasoning capabilities (such as GPT-o3*, Deepseek-R1*) perform better in command understanding tasks such as long text understanding and multi-round dialogue, but have not completely solved all problems. In the intention reasoning task, except for "Inconsistent information reasoning", the models generally performed poorly, and only a few models partially passed the case. In terms of reliable dialogue generation, all models performed poorly in response stability, and only GPT-o3 and Deepseek-R1 performed well in the alignment test. Overall, while reasoning abilities contribute to improved task performance, significant limitations remain across all three stages.
@@ 6 Benchmark
This section covers benchmarks for LLMs in above three stage (Table 1), more details see Appendix A.
@@ 6.1 Benchmarking Instruction Understanding
Instruction understanding in LLMs involves extracting key information, maintaining coherence, and adapting to dynamic conversation changes, especially in long or multi-round dialogues. LLM capabilities in multi-turn dialogue and long-context processing have been explored through various benchmarks. BotChat Duan et al. (2024) evaluates dialogue generation, showing GPT-4’s strengths but noting instruction compliance and length limitations in other models. MINT Wang et al. (2024) highlights limited progress in tool use and feedback for complex tasks. MT-Bench-101 Bai et al. (2024) identifies challenges in enhancing long-term interaction skills. For long-context tasks, performance drops in ultra-long texts (Zhang et al. Zhang et al. (2024c)), L-Eval An et al. (2024) emphasizes length-instruction-enhanced metrics, and LongICLBench Li et al. (2025) reveals difficulties in reasoning across multiple pieces of information.
Stage Challenge GPT-4o GPT-o3* Qwen3 Qwen3* Deepseek-v3 Deepseek-R1*
Instruction Understanding Long-Text Comprehension C A C A C A
Multi-Turn Conversation Handling C A C C C A
Intention Reasoning Inconsistent Instruction Reasoning C C C C B C
Misinformation Reasoning B A C B B A
Fuzzy Language Interpretation B B C C C C
Intention Clarification Failure C B C C C C
Reliable Dialog Generation Response Stability C C C C C C
Alignment C A C C C A
Table 2: Results of using different LLMs on the challenge cases; * denotes models with reasoning abilities. ’A’ indicates the model’s accuracy is greater than 75%, ’B’ is between 50% and 75%, and ’C’ is below 50%.
@@ 6.2 Benchmarking LLM Reasoning
LLM intention reasoning involves inferring user intentions by interpreting both explicit and implicit language cues. The existing benchmarks comprehensively evaluate the multifaceted reasoning capabilities of LLMs, encompassing vulnerability to indirect hint injection attacks (BIPIA Yi et al. (2025)), advantages in multimodal intention understanding (Miko Lu et al. (2024)), analysis of self-contradictions in long documents (CONTRADOC Li et al. (2024a)), and contradiction detection in Chinese dialogues (CDCONV Zheng et al. (2022)). Collectively, these benchmarks highlight both the challenges and advancements of LLMs in complex reasoning.
@@ 6.3 Benchmarking LLM Generation
LLM Generation assesses a model’s ability to understand user instructions, avoid fabricating false information, and generate accurate, contextually appropriate responses. Open-LLM-Leaderboard Ye et al. (2024), ETHICS Hendrycks et al. (2021) and FACTOR Muhlgay et al. (2024) all focus on the reliability evaluation of content generated by large models. Open-LLM-Leaderboard finds that large-scale models have higher uncertainty, and fine-tuned models have higher accuracy but greater uncertainty. ETHICS focuses on the ethical value alignment of generated content. FACTOR evaluates factuality through scalable methods to ensure that diverse and rare facts are covered.
@@ 7 Future Directions
This section summarizes ongoing challenges in above stages with LLMs and outlines potential future research directions.
@@ 8 Conclusion
This paper analyzes LLMs’ performance in processing user instructions. Despite progress in natural language understanding, LLMs struggle with complex, inconsistent instructions, often resulting in biases, errors, and hallucinations. Improvements through prompt engineering, model expansion, and RLHF et al. have not fully addressed LLMs’ limitations in reasoning and comprehension, limiting their real-world applicability. We identify three challenges: instruction understanding, intention reasoning and reliable dialog generation. Future research should focus on enhancing reasoning for complex instructions and aligning outputs with user intention to improve LLMs’ reliability.
Limitations
We have made significant efforts to ensure the quality of this study, but certain limitations may still exist. First, due to constraints on the length of the discussion, this article primarily focuses on analyzing examples of human interaction with large models, which inevitably limits the depth of methodological details we can explore. Second, while our meta-analysis draws extensively from prominent academic platforms such as the ACL conference series, ICLR, ICML, and the arXiv preprint repository, it is possible that some valuable insights from other sources have not been included. It is worth noting that several open scientific questions in this field remain unresolved, and the academic community has yet to reach a consensus on these issues. To address these limitations, we plan to establish a long-term tracking mechanism to monitor new developments in the field. This will allow us to incorporate emerging theoretical advancements and dynamically refine or expand upon the perspectives presented in this study.
Appendix A Benchmark Details
A.1 Instruction Understanding Benchmark
BotChat Duan et al. (2024) specifically addresses the evaluation of LLMs’ ability to emulate human-like, multi-turn conversations using an LLM-centric approach. This benchmark assesses utterance generation, evaluation protocols, and judgment. Findings suggest that models like GPT-4 demonstrate exceptional performance as both generators and judges, producing human-indistinguishable dialogues and showing high alignment with human evaluation standards. Conversely, other LLMs face challenges in generating quality multi-turn dialogues due to issues like inadequate instruction-following and a tendency towards prolixity, particularly in generating extensive dialogues.
MINT Wang et al. (2024) introduces a benchmark that evaluates LLMs’ capacity to solve complex tasks through multi-turn interactions, emphasizing the use of external tools and leveraging natural language feedback. MINT provides a reproducible evaluation framework where LLMs can access tools and receive simulated user feedback. Analysis using MINT reveals that LLMs generally benefit from tools and language feedback. Interestingly, studies with MINT suggest that techniques like supervised instruction-finetuning (SIFT) and reinforcement learning from human feedback (RLHF) might not consistently improve multi-turn capabilities, indicating the need for further research in training methodologies for conversational settings.
MT-BENCH-101 Bai et al. (2024) offers a structured approach with its three-layer hierarchical capability taxonomy and a dataset of 1,388 dialogue pairs across 13 tasks. This allows for a fine-grained evaluation of LLMs’ multi-turn dialogue skills. The findings from studies utilizing MT-BENCH-101 indicate that commonly employed techniques yield limited improvements in multi-turn performance, suggesting a need for more effective methods to enhance conversational abilities over extended interactions.
Zhang et al. Zhang et al. (2024c) focus on the challenge of long-context processing, introducing a benchmark of 12 tasks, each averaging over 100K tokens, to assess LLMs’ long-context processing abilities. Results indicate a marked performance drop with longer contexts, highlighting the need for further improvement.
L-Eval An et al. (2024) contributes to standardizing the evaluation of Long-Context Language Models (LCLMs) by providing a new evaluation suite with diverse tasks, document lengths (3k to 200k tokens), and a large number of human-labeled query-response pairs. It also investigates the effectiveness of evaluation metrics, suggesting that Length-instruction-enhanced (LIE) evaluation and LLM judges correlate better with human judgments. Comprehensive studies using L-Eval offer insights into the performance of both commercial and open-source LCLMs.
LongICLBench Li et al. (2025) focuses on long in-context learning in extreme-label classification, with input lengths ranging from 2K to 50K tokens and a large number of classes. Evaluation on this benchmark showed that while LLMs perform well on less challenging tasks, they struggle with more complex ones involving a large label space and longer contexts, revealing a bias towards later parts of the input and difficulty in reasoning over multiple pieces of information.
A.2 LLM Intention Reasoning Benchmark
BIPIA Yi et al. (2025) evaluates LLMs under indirect hint injection attacks across five scenarios and 250 targets, revealing vulnerabilities in all models, with GPT-3.5-turbo and GPT-4 exhibiting notably higher susceptibilities.
Miko Lu et al. (2024) assesses multimodal models in understanding social media user intentions. The benchmark, which includes 979 social media entries, shows that multimodal LLMs outperform text-only models like LLama2-7B and GLM4. Incorporating image data enhances the model’s ability to interpret user intentions, improving accuracy.
CONTRADOC Li et al. (2024a) is the first dataset for analyzing self-contradictions in long documents. Evaluation of GPT-4, PaLM2, and other LLMs on this dataset reveals that while GPT-4 outperforms others and even surpasses human performance, it still struggles with complex contradictions requiring nuanced reasoning.
CDCONV Zheng et al. (2022) focuses on contradiction detection in Chinese dialogues, containing over 12,000 dialogue rounds. The study shows that the Hierarchical method consistently outperforms others in detecting contradictions, highlighting the importance of accurate contextual modeling in dialogue understanding.
A.3 LLM Reliable Generation Benchmark
Open-LLM-Leaderboard Ye et al. (2024) introduces a novel benchmark that integrates uncertainty quantification to evaluate the reliability of content generation across tasks like QA, comprehension, and dialogue. Results show that larger models often exhibit greater uncertainty, and fine-tuned models tend to have higher uncertainty despite higher accuracy.
ETHICS Hendrycks et al. (2021) evaluates whether generated content aligns with human ethical values, such as justice and well-being. The study finds that while models like GPT-3 show promise in predicting human moral judgments, they still need improvement in this domain.
FACTOR Muhlgay et al. (2024) addresses the evaluation of factuality in LLMs by providing a scalable method that ensures diverse and rare facts are considered. Testing models such as GPT-2 and GPT-Neo show that, while benchmark scores correlate with perplexity, they better reflect factuality in open-ended generation, especially when retrieval augmentation is applied.
Appendix B Challenges Cases in Human-LLMs Aligement
To fully understand the limitations of large language models (LLMs) in practical applications, it is particularly important to analyze their performance in a variety of complex scenarios. We collected human user instructions from real scenarios that the large model needs to interact with on a total of eight challenges that the summarized existing large language models (LLMs) face in the three phases of instruction understanding, intention reasoning, and reliable dialog generation, and cleaned and filtered the instructions through manual screening, diversity sampling, and difficulty filtering, and finally, 50 human instructions were used for each challenge.
The statistical results of the test are shown in Table.2. In this section, we will present a series of representative case studies and conduct an in-depth analysis of the current mainstream LLMs (including GPT-4o, GPT-o3, Qwen3, DeepSeek-V3 and DeepSeek-R1). Besides, the blue smiley face indicates that the model is capable of providing accurate responses or identifying inconsistencies in the user’s input instructions. In contrast, the red crying face signifies that the LLM failed to recognize contradictions in the user’s instructions and produced incorrect responses. Through these cases, we systematically reveal the typical failure modes of each model in different aspects and their deep-seated causes, providing important references and directions for targeted optimization in future research.
B.1 Case of Instruction Understanding
Although large language models (LLMs) perform well in short text and single-round dialogue scenarios, their ability to understand and execute complex instructions in long contexts and multi-round dialogues still faces many severe challenges.
Figure 4 shows that long texts usually contain a lot of irrelevant or redundant information, which causes the core content directly related to the task to be submerged. The model is prone to omission or confusion when extracting key information, and even hallucinations, thereby generating content that is irrelevant to the user’s instructions.
Figure 5 shows that when dealing with long-distance information associations that need to span multiple paragraphs or dialogue rounds, the model often has difficulty tracking and integrating the logical relationship between contexts, thereby losing important clues. In multi-round dialogue scenarios, the model is not only prone to gradually accumulate errors due to inaccurate understanding of the previous text, but may also fail to correctly judge the relevance of each round of dialogue, causing the answer to deviate from the user’s real needs.
As shown in Table 2, Figure 4, and Figure 5, in the instruction understanding stage, the reasoning models generally performed well on the "Long-Text Comprehension" task, while the non-reasoning models all struggled, and for “Multi-Turn Conversation Handling”, in addition to the non-reasoning models, Qwen3, which provides reasoning capability, also exhibited suboptimal performance. The above results show that reasoning can effectively improve the ability of LLMs in instruction understanding, but cannot completely solve the instruction understanding problem.
B.2 Case of Intention Reasoning
Since there are often spelling errors, factual contradictions and semantic ambiguities in user instructions, LLM faces many challenges in understanding the user’s true intentions.
Figure 6 shows that the model often relies too much on the user’s input and ignores the obvious errors in the input. It tends to generate answers directly instead of identifying and correcting these errors first;
Figure 7 shows that when knowledge updates are not synchronized or input data is maliciously tampered with, the model is more likely to output information that is inconsistent with actual needs or contaminated. However, the reasoning model can identify the errors, indicating that the model’s reasoning ability is crucial in identifying the user’s input intention.
Figure 8 shows that when faced with ambiguous or uncertain expressions, large models usually tend to customize questions or give default explanations based on their own training preferences, rather than actively asking users for more context.The GPT series performs better than models such as DeepSeek in such tasks, indicating that even models with strong reasoning capabilities may have difficulties in dealing with modal expressions such as sarcasm and irony. This reflects the limitations of AI in understanding complex human language expressions, as well as differences in the coverage of such language phenomena in the training data and the depth of the model’s understanding of the social and cultural context.
Figure 9 shows that for implicit intentions such as sarcasm, metaphors, and emotions, the model often only focuses on the literal meaning and has difficulty grasping the deep emotions or context, thereby outputting incorrect analysis results.
As shown in Table 2, Figure 6 to Figure 9, in the intention reasoning stage, all models encountered difficulties and performed poorly in “Inconsistent Instruction Reasoning”, “Fuzzy Language Interpretation” and “Intention Clarification Failure”; in “Misinformation Reasoning”, only GPT-o3 and Deepseek-R1 achieved good performance, while the other models underperformed. This suggests that all models have significant challenges in reasoning about intentions.
B.3 Case of Reliable Dialog Generation
Current large language models exhibit both deterministic response preferences and ambiguous knowledge boundaries during generation. When queries fall within the model’s knowledge coverage, the outputs are generally reliable; however, in open-domain or previously unseen scenarios, the quality of generated content fluctuates significantly.
Figure 10 illustrates the instability of LLM-generated content, which primarily manifests in two typical issues: First, when the model lacks relevant knowledge or information regarding the input (Q.1), it tends to fabricate details, producing content inconsistent with facts and thereby substantially compromising the accuracy and reliability of its outputs. Second, even when the model possesses relevant knowledge (Q.2), it may still misinterpret or improperly integrate contextual information, resulting in outputs that are contextually inappropriate or logically flawed. These issues not only undermine user experience but also limit the applicability of LLMs in high-stakes scenarios.
Figure 11 demonstrates the misalignment between model-generated content and human values, mainly reflected in the model’s potential to violate widely accepted moral and ethical standards. On one hand, when handling sensitive or controversial topics, the model may generate harmful, offensive, or discriminatory statements, negatively impacting users. On the other hand, it may also produce responses that encourage illegal activities or contravene social ethics, introducing potential legal and societal risks and imposing higher requirements on the safety and trustworthiness of LLMs.
As shown in Table 2, Figure 10 and Figure 11, in the reliable dialog generation stage, all models failed to maintain “Response Stability”; in “Alignment” test, only GPT-o3 and Deepseek-R1 performed well, while the rest of the models failed. The results show that all models have clear shortcomings in “Response Stability”, and the reasoning models have improved in “Alignment”.
Overall, reasoning capabilities can improve the performance of llms for interacting with real human user commands, however, all models still face significant challenges and remain underpowered in real-world and human interaction scenarios.
Figure 4: Case of Remote Information Failure (§2.1), where the model forgets relevant information over long distances in long context.
Figure 5: Case of Incorrect Relevance Judgment2.2), such as the model incorrectly associates wrong content from the previous turn.
Figure 6: Case of Inconsistent Instruction Reasoning (§3.1), where LLMs fail to detect conflicting inputs (Q.1) or overlooking logical inconsistencies (Q.2)
Figure 7: Case of Misinformation Reasoning3.2), caused by temporal misalignment leading to failure responses (Q.1) or data contamination resulting in misleading outputs (Q.2).
Figure 8: Case of Fuzzy Language Interpretation3.3), where the model relies on biases for fuzzy queries (C.1) or defaults to a response without seeking clarification (C.2)
Figure 9: Case of Intention Clarification Failure3.4) , where it misinterprets sarcasm (Q.1) or ignores prior emotional context (Q.2)
Figure 10: Case of Unstable Content Generation (§4.1), where the model fabricates details when it lacks relevant knowledge (Q.1) or produces incorrect contextual information despite having relevant knowledge (Q.2)
Figure 11: Case of Misalignment with Human Values (§4.2), where the model generates harmful or offensive content (Q.1)